There were 3 girls ( T , K, G). G gave half of her sweets to T. T gave half of her lollies to K, then K gave G half of her lollies. If G had the same amount of lollies at the start, how much lollies did G have?
Please help me to set up equations to solve this problem. Thanks!

Answer by stanbon(75887)

There were 3 girls ( T , K, G).
----G gave half of her sweets to T.
G now has G/2
T now has T+(G/2) = (2T+G)/2
--------------------------------------
T gave half of her lollies to K
T now has (2T+G)/4
K now has K + (2T+G)/4 = (4K+2T+G)/8
---------------
then K gave G half of her lollies
K now has (4K+2T+G)/16
G now has (G/2)+(4k+2T+G)/16 = (8G+4K+2T+G)/16 = (9G + 4K+2T)/16
=========
If G had the same amount of lollies at the start, how much lollies did G have?
Please help me to set up equations to solve this problem.
Solve for "G"::
(9G+4K+2T)/16 = G
(9G + 4K + 2T) = 16G
7G = 4K + 2T
G = (4K+2T)/7
----------------
Cheers,
Stan H.
----------------